# Towards Generalizable PDE Dynamics Forecasting via Physics-Guided Invariant Learning

### v1

@@ Abstract
Advanced deep learning-based approaches have been actively applied to forecast the spatiotemporal physical dynamics governed by partial differential equations (PDEs), which acts as a critical procedure in tackling many science and engineering problems. As real-world physical environments like PDE system parameters are always capricious, how to generalize across unseen out-of-distribution (OOD) forecasting scenarios using limited training data is of great importance. To bridge this barrier, existing methods focus on discovering domain-generalizable representations across various PDE dynamics trajectories. However, their zero-shot OOD generalization capability remains deficient, since extra test-time samples for domain-specific adaptation are still required. This is because the fundamental physical invariance in PDE dynamical systems are yet to be investigated or integrated. To this end, we first explicitly define a two-fold PDE invariance principle, which points out that ingredient operators and their composition relationships remain invariant across different domains and PDE system evolution. Next, to capture this two-fold PDE invariance, we propose a physics-guided invariant learning method termed iMOOE, featuring an Invariance-aligned Mixture Of Operator Expert architecture and a frequency-enriched invariant learning objective. Extensive experiments across simulated benchmarks and real-world applications validate iMOOE’s superior in-distribution performance and zero-shot generalization capabilities on diverse OOD forecasting scenarios.
@@ 2.1 Problem Formulation
where $\mathbf{I}^{e}=\{\mathbf{u}^{e}(t,\mathbf{x})\}_{t=0}^{H-1}$ is a trajectory of past $H$ steps observed from $e$, $\mathbf{Y}^{e}=\{\mathbf{u}^{e}(t,\mathbf{x})\}_{t=H}^{N_{t}}$ is the target sequence that should be predicted. $\ell(\cdot)$ is the loss function quantifying prediction errors. For brevity, we use $\mathbf{u}_{t}$ to denote $\mathbf{u}(t,\mathbf{x})$ in the rest of content. We describe the practical value of such zero-shot OOD dynamics forecasting setting in Appendix B.4.
@@ 2.2 Invariant Learning for Dynamics Forecasting
Figure 1: (a) The SCM diagram for the formation process PDE dynamics. It illustrates prescribed two-level PDE invariance and potential distribution shifts on exogenous inputs. (b) A case study by varying physical parameters of DR dynamics to compare the zero-shot OOD performance of four methods. Without the guidance of formalized PDE invariance, previous methods can not achieve better OOD results on unseen environments. (c) The ID-OOD correlations of two neural operators. Based on the slope of two linear positive ID-OOD lines, FNO equipped with PDE invariance learning can capture more transferrable knowledge from limited training domains and achieve better OOD robustness. Refer to Appendix E.6 for more details.
@@ 2.3 Invariance Principle for PDE Dynamics
In addition, the future state $\mathbf{\hat{u}}_{t+1}$ can be calculated as $\mathbf{\hat{u}}_{t+1}=\int_{t}^{t+1}h(\{\sigma_{i}\}_{i=1}^{K},\mathbf{p},\mathbf{f})\mathrm{d}t+\mathbf{u}_{t}$ given the last observation $\mathbf{u}_{t}$. In a nutshell, invariant correlations between input $\mathbf{I}$ and target $\mathbf{Y}$ involve: invariant operators $\{\sigma_{i}\}_{i=1}^{K}$, invariant compositional relationships $h$ among different items, as well as the fixed step-wise numerical integration. We present a structural causal model (SCM) (Krueger et al., 2021) in Fig. 1(a) to illustrate the formation process of PDE dynamics and its pertinent two-level invariant correlations. Regardless of various time steps and distribution shifts on $\{\mathbf{u}_{0},\mathbf{p},\mathbf{f}\}$, the fundamental set of operators and their composition relationships can remain invariant. Besides, we provide a case study by varying diffusion and reaction coefficient of DR dynamics in Fig. 1(b), and the ID-OOD correlation lines (an effective metric to assess OOD robustness (Yuan et al., 2023)) in Fig. 1(c). Both of them can further demonstrate the effectiveness of the proposed physics-guided PDE invariance learning for improving zero-shot OOD capability. See Appendix B for more related works on PDE dynamics forecasting and invariant learning.
@@ 3.1 Architecture Alignment: Mixture of Operator Experts
We illustrate this masking-based input derivative selection design in Fig. 2(b), and analyze its effect in Appendix E.1.2.
@@
In Appendix E.1.3, we verify that choosing a suitable type of fusion network can better align with the compositionality invariance and achieve better OOD performance.
@@ 3.2 Frequency-Enriched Invariant Learning Objective
where $\mathcal{R}^{e}_{pred}$ is provided in Eq. 6. We borrow the useful linear scheduling scheme to impose this risk equality loss (Krueger et al., 2021), which reserves an initial empirical risk minimization stage (i.e. a pretraining stage merely by $\mathcal{L}_{pred}$) to learn the rich predictive representations. Refer to Appendix E.1.5 for the effect of this linear invariant loss scheduling. Note that our design differs from (Krueger et al., 2021) in environment division. In addition to taking physical parameters in $\mathbf{p}$ and $\mathbf{f}$ as environment labels, we also partition training domains by different autoregressive steps, since there exist covariate shifts in $p(\mathbf{I}^{e},\mathbf{Y}^{e})$. Specifically, during the autoregressive prediction, the distribution of past sequences $p(\mathbf{I}^{e})$ can change with the time marching, but the correlations between $\mathbf{I}^{e}$ and $\mathbf{Y}^{e}$ keep invariant at each time step. We find this step-wise division is instrumental for fluid dynamics forecasting such as Navier-Stokes and Burgers systems, as shown in Appendix E.1.4.
@@
where $\mathcal{F}$ is the fast Fourier transform and $\xi$ is the wavenumber vector for each spatial frequency. Apparently, the weight $\xi||_{2}^{2}$ can pay more attention to the high-frequency modes at each forecasting step. We validate such frequency enrichment loss can induce better PDE forecasting generalization in Appendix. E.1.1. Prior works on OOD vision recognition (Chen et al., 2023a; Zhang et al., 2022) also proved that diverse and rich features can lead to better OOD capability.
@@ 3.3 Overall Framework
Equipped with this hybrid training loss and invariance-aligned architecture developed in Sec. 3.1, we can effectively learn the proposed PDE invariance to achieve zero-shot OOD forecasting. Existing neural operators always train with prediction loss $\mathcal{L}_{pred}$, without any effort to learn the fundamental PDE invariance principle. This could be the key reason for their failures on OOD dynamics forecasting. We demonstrate in Section 4.3 that when equipped with the explicit physics-informed PDE invariance learning method iMOOE, current neural operators can realize better OOD performance. Moreover, in Appendix E.7, we further investigate how the properties of multi-environment training data can affect the zero-shot OOD capability of iMOOE. As simulating PDE trajectories or measuring real-world PDE dynamics is expensive, such analysis can provide a guideline on how to collect training data under a limited data budget.
@@ 4.1 Experimental Setup
Datasets. We adopt five PDE dynamical systems in different fields for the spatiotemporal physical dynamics forecasting task: Diffusion-Reaction (DR) (Takamoto et al., 2022), Navier-Stokes (NS) (Li et al., 2021), Burgers (BG) (Hao et al., 2024b), Shallow-Water (Takamoto et al., 2022) and Heat-Conduction (HC) (Hao et al., 2024b). We construct a wide range of OOD scenarios by varying the physical parameters of initial conditions $\mathbf{u}_{0}$, PDE coefficients $\mathbf{p}$, forcing terms $\mathbf{f}$ or temporal resolutions $N_{t}$. ID and OOD parameters for PDE simulation are randomly drawn from two non-overlapped uniform distributions, while previous parametric PDE learning works like (Kassaï Koupaï et al., 2024; Takamoto et al., 2023) just select several separate parameters. The spatial resolution of each state frame is fixed to $64\times 64$. See Appendix C for detailed description on data generation.
@@ 4.2 Zero-shot OOD Performance
Baselines. We select six latest PDE forecasting methods with highlighted OOD generalization capability: i) CoDA (Kirchmeyer et al., 2022) and GEPS (Kassaï Koupaï et al., 2024): two context-aware meta-learning-based models. ii) CAPE (Takamoto et al., 2023): a parameter conditioning method. iii) CNO (Raonic et al., 2023): a robust convolutional neural operator. iv) DPOT (Hao et al., 2024a): a transformer-based operator with denoising pretraining. v) VCNeF (Hagnberger et al., 2024): a conditional neural field-based method. Note that meta-learning-based methods commonly require few-shot adaptation to perform OOD forecasting. In Appendix E.4, we describe how to adapt them to zero-shot setting and further compare zero-shot iMOOE with few-shot CoDA, GEPS. Implementation details of these baseline models are provided in Appendix E.9.
@@
Results. We report ID/OOD generalization outcomes on various unseen scenarios in Table 1. It is obvious that iMOOE can achieve the state-of-art (SOTA) results on this simulated benchmark, with an average increase of $40.21\%$ on nMSE and $30.78\%$ on fRMSE. Such considerable promotion reflects that explicitly learning the proposed physics-guided PDE invariance can boost zero-shot OOD performance on PDE dynamics forecasting. Moreover, we present the OOD results on extrapolated temporal resolutions in Table 2. Following previous time extrapolation setting (Kassaï Koupaï et al., 2024), we train on $0,N_{t}]$ and test on $0,2N_{t}]$. We find that iMOOE can achieve SOTA results with an average growth of $32.51\%$ on nMSE and $15.30\%$ on fRMSE. It indicates that learning the underlying PDE invariance across time steps can improve the OOD performance on unseen temporal distribution shift scenarios. To measure iMOOE’s zero-shot OOD capacity more clearly, we present an empirical upper bound for its OOD performance in Appendix E.5.
@@ 4.4 Application to Real-world PDE Dynamics
Apart from the simulated benchmark above, we also leverage real-world PDE-governed dynamics of daily Sea Surface Temperature (SST) (Huang et al., 2021) for OOD validation. A specific region on Pacific Ocean is selected and divided into $60\times 60$ grid. We predict SST state of future $6$ days using past $4$ days’ observations. SST data between year 19821982-$2019$ and 20202020-$2021$ is utilized for training and testing. Note that each independent SST trajectory can be deemed as an instance from a unique environment, since daily SST variations are affected by many exogenous conditions like solar radiation and ocean currents. Input parameters are also unknown so we feed one-valued vector to the fusion network. We take a typical SST baseline called DyAd (Wang et al., 2022b) and present the comparison results in Table 3. We can see that iMOOE can attain the lowest mean and variance of two metrics across various OOD samples. It reflects that iMOOE can better estimate the invariant physical laws in SST. Test samples in Fig. 3 exhibit iMOOE can capture local variations of SST with higher fidelity and accuracy. Beyond 2D PDE-governed dynamics, we further demonstrate iMOOE’s zero-shot OOD capability can be extended to other types of dynamical systems in Appendix E.2.
4.5 Sensitivity analysis
@@
Number of
expert $K$
DR BG
Inference
time
@@
In Table 5, we investigate the influence of the number of operator experts $K$ on iMOOE’s zero-shot OOD performance. To ensure a fair comparison, we only escalate $K$ from 1 to 4 and keep other setups like the width of FNO and training batch size unchanged. We find that the best-performing group is $K=3$ while the worst setting is $K=1$. This reveals that small $K$ (i.e. only 1 expert) is not sufficient to capture the operator invariance, while large $K$ (i.e. 4 experts) could be redundant given that actual PDE systems contains only a few number of compositional invariant operators (Rudy et al., 2017). Besides, the increasing number of neural operators can exacerbate the computational overhead. Refer to Appendix D for more detailed explanations on the effect of expert number $K$ and its distinction from the mixture-of-expert (MoE) architecture in large foundation models (LFMs). Refer to Appendix E.3 for more sensitivity analysis on loss weights in Eq. 9.

### v2

# arXiv:2509.24332v2[cs.LG] 28 Feb 2026

@@ Abstract
Advanced deep learning-based approaches have been actively applied to forecast the spatiotemporal physical dynamics governed by partial differential equations (PDEs), which acts as a critical procedure in tackling many science and engineering problems. As real-world physical environments like PDE system parameters are always capricious, how to generalize across unseen out-of-distribution (OOD) forecasting scenarios using limited training data is of great importance. To bridge this barrier, existing methods focus on discovering domain-generalizable representations across various PDE dynamics trajectories. However, their zero-shot OOD generalization capability remains deficient, since extra test-time samples for domain-specific adaptation are still required. This is because the fundamental physical invariance in PDE dynamical systems are yet to be investigated or integrated. To this end, we first explicitly define a two-fold PDE invariance principle, which points out that ingredient operators and their composition relationships remain invariant across different domains and PDE system evolution. Next, to capture this two-fold PDE invariance, we propose a physics-guided invariant learning method termed iMOOE, featuring an Invariance-aligned Mixture Of Operator Expert architecture and a frequency-enriched invariant learning objective. Extensive experiments across simulated benchmarks and real-world applications validate iMOOE’s superior in-distribution performance and zero-shot generalization capabilities on diverse OOD forecasting scenarios. The code is be publicly accessible at https://github.com/LSY-Cython/iMOOE.
@@ 2.1 Problem Formulation
where $\mathbf{I}^{e}=\{\mathbf{u}^{e}(t,\mathbf{x})\}_{t=0}^{H-1}$ is a trajectory of past $H$ steps observed from $e$, $\mathbf{Y}^{e}=\{\mathbf{u}^{e}(t,\mathbf{x})\}_{t=H}^{N_{t}}$ is the target sequence that should be predicted. $\ell(\cdot)$ is the loss function quantifying prediction errors. For brevity, we use $\mathbf{u}_{t}$ to denote $\mathbf{u}(t,\mathbf{x})$ in the rest of content. We describe the practical value of such zero-shot OOD dynamics forecasting setting in Appendix C.4.
@@ 2.2 Invariant Learning for Dynamics Forecasting
Figure 1: (a) The SCM diagram for the formation process PDE dynamics. It illustrates prescribed two-level PDE invariance and potential distribution shifts on exogenous inputs. (b) A case study by varying physical parameters of DR dynamics to compare the zero-shot OOD performance of four methods. Without the guidance of formalized PDE invariance, previous methods can not achieve better OOD results on unseen environments. (c) The ID-OOD correlations of two neural operators. Based on the slope of two linear positive ID-OOD lines, FNO equipped with PDE invariance learning can capture more transferrable knowledge from limited training domains and achieve better OOD robustness. Refer to Appendix F.6 for more details.
@@ 2.3 Invariance Principle for PDE Dynamics
In addition, the future state $\mathbf{\hat{u}}_{t+1}$ can be calculated as $\mathbf{\hat{u}}_{t+1}=\int_{t}^{t+1}h(\{\sigma_{i}\}_{i=1}^{K},\mathbf{p},\mathbf{f})\mathrm{d}t+\mathbf{u}_{t}$ given the last observation $\mathbf{u}_{t}$. In a nutshell, invariant correlations between input $\mathbf{I}$ and target $\mathbf{Y}$ involve: invariant operators $\{\sigma_{i}\}_{i=1}^{K}$, invariant compositional relationships $h$ among different items, as well as the fixed step-wise numerical integration. We present a structural causal model (SCM) (Krueger et al., 2021) in Fig. 1(a) to illustrate the formation process of PDE dynamics and its pertinent two-level invariant correlations. Regardless of various time steps and distribution shifts on $\{\mathbf{u}_{0},\mathbf{p},\mathbf{f}\}$, the fundamental set of operators and their composition relationships can remain invariant. We formally justify the proposed two-level invariance principle can hold across many specialized PDE systems from three perspectives in Appendix B. Besides, we provide a case study by varying diffusion and reaction coefficient of DR dynamics in Fig. 1(b), and the ID-OOD correlation lines (an effective metric to assess OOD robustness (Yuan et al., 2023)) in Fig. 1(c). Both of them can further demonstrate the effectiveness of the proposed physics-guided PDE invariance learning for improving zero-shot OOD capability. Refer to Appendix C for more related works on PDE dynamics forecasting and invariant learning.
@@ 3.1 Architecture Alignment: Mixture of Operator Experts
We illustrate this masking-based input derivative selection design in Fig. 2(b), and analyze its effect in Appendix F.1.2.
@@
In Appendix F.1.3, we verify that choosing a suitable type of fusion network can better align with the compositionality invariance and achieve better OOD performance.
@@ 3.2 Frequency-Enriched Invariant Learning Objective
where $\mathcal{R}^{e}_{pred}$ is provided in Eq. 6. We borrow the useful linear scheduling scheme to impose this risk equality loss (Krueger et al., 2021), which reserves an initial empirical risk minimization stage (i.e. a pretraining stage merely by $\mathcal{L}_{pred}$) to learn the rich predictive representations. Refer to Appendix F.1.5 for the effect of this linear invariant loss scheduling. Note that our design differs from (Krueger et al., 2021) in environment division. In addition to taking physical parameters in $\mathbf{p}$ and $\mathbf{f}$ as environment labels, we also partition training domains by different autoregressive steps, since there exist covariate shifts in $p(\mathbf{I}^{e},\mathbf{Y}^{e})$. Specifically, during the autoregressive prediction, the distribution of past sequences $p(\mathbf{I}^{e})$ can change with the time marching, but the correlations between $\mathbf{I}^{e}$ and $\mathbf{Y}^{e}$ keep invariant at each time step. We find this step-wise division is instrumental for fluid dynamics forecasting such as Navier-Stokes and Burgers systems, as shown in Appendix F.1.4.
@@
where $\mathcal{F}$ is the fast Fourier transform and $\xi$ is the wavenumber vector for each spatial frequency. Apparently, the weight $\xi||_{2}^{2}$ can pay more attention to the high-frequency modes at each forecasting step. We validate such frequency enrichment loss can induce better PDE forecasting generalization in Appendix. F.1.1. Prior works on OOD vision recognition (Chen et al., 2023a; Zhang et al., 2022) also proved that diverse and rich features can lead to better OOD capability.
@@ 3.3 Overall Framework
Equipped with this hybrid training loss and invariance-aligned architecture developed in Sec. 3.1, we can effectively learn the proposed PDE invariance to achieve zero-shot OOD forecasting. Existing neural operators always train with prediction loss $\mathcal{L}_{pred}$, without any effort to learn the fundamental PDE invariance principle. This could be the key reason for their failures on OOD dynamics forecasting. We demonstrate in Section 4.3 that when equipped with the explicit physics-informed PDE invariance learning method iMOOE, current neural operators can realize better OOD performance. Moreover, in Appendix F.7, we further investigate how the properties of multi-environment training data can affect the zero-shot OOD capability of iMOOE. As simulating PDE trajectories or measuring real-world PDE dynamics is expensive, such analysis can provide a guideline on how to collect training data under a limited data budget.
@@ 4.1 Experimental Setup
Datasets. We adopt five PDE dynamical systems in different fields for the spatiotemporal physical dynamics forecasting task: Diffusion-Reaction (DR) (Takamoto et al., 2022), Navier-Stokes (NS) (Li et al., 2021), Burgers (BG) (Hao et al., 2024b), Shallow-Water (Takamoto et al., 2022) and Heat-Conduction (HC) (Hao et al., 2024b). We construct a wide range of OOD scenarios by varying the physical parameters of initial conditions $\mathbf{u}_{0}$, PDE coefficients $\mathbf{p}$, forcing terms $\mathbf{f}$ or temporal resolutions $N_{t}$. ID and OOD parameters for PDE simulation are randomly drawn from two non-overlapped uniform distributions, while previous parametric PDE learning works like (Kassaï Koupaï et al., 2024; Takamoto et al., 2023) just select several separate parameters. The spatial resolution of each state frame is fixed to $64\times 64$. See Appendix D for detailed description on data generation.
@@ 4.2 Zero-shot OOD Performance
Baselines. We select six latest PDE forecasting methods with highlighted OOD generalization capability: i) CoDA (Kirchmeyer et al., 2022) and GEPS (Kassaï Koupaï et al., 2024): two context-aware meta-learning-based models. ii) CAPE (Takamoto et al., 2023): a parameter conditioning method. iii) CNO (Raonic et al., 2023): a robust convolutional neural operator. iv) DPOT (Hao et al., 2024a): a transformer-based operator with denoising pretraining. v) VCNeF (Hagnberger et al., 2024): a conditional neural field-based method. Note that meta-learning-based methods commonly require few-shot adaptation to perform OOD forecasting. In Appendix F.4, we describe how to adapt them to zero-shot setting and further compare zero-shot iMOOE with few-shot CoDA, GEPS. Implementation details of these baseline models are provided in Appendix F.9.
@@
Results. We report ID/OOD generalization outcomes on various unseen scenarios in Table 1. It is obvious that iMOOE can achieve the state-of-art (SOTA) results on this simulated benchmark, with an average increase of $40.21\%$ on nMSE and $30.78\%$ on fRMSE. Such considerable promotion reflects that explicitly learning the proposed physics-guided PDE invariance can boost zero-shot OOD performance on PDE dynamics forecasting. Moreover, we present the OOD results on extrapolated temporal resolutions in Table 2. Following previous time extrapolation setting (Kassaï Koupaï et al., 2024), we train on $0,N_{t}]$ and test on $0,2N_{t}]$. We find that iMOOE can achieve SOTA results with an average growth of $32.51\%$ on nMSE and $15.30\%$ on fRMSE. It indicates that learning the underlying PDE invariance across time steps can improve the OOD performance on unseen temporal distribution shift scenarios. To measure iMOOE’s zero-shot OOD capacity more clearly, we present an empirical upper bound for its OOD performance in Appendix F.5.
@@ 4.3 Universality Study
Table 5: Zero-shot OOD results on SSE dynamics.
Models nMSE fRMSE
Mean Std Mean Std
GEPS 3.41e-2 4.96e-3 4.11e-3 1.55e-4
[…]
VCNeF 3.41e-2 4.96e-3 4.11e-3 1.55e-4
DyAd 3.17e-2 4.61e-3 3.96e-3 1.50e-4
CAPE 3.34e-2 4.85e-3 4.07e-3 1.53e-4
iMOOE 1.52e-2 2.34e-3 2.78e-3 1.42e-4
Figure 4: Test SSE sample showcase.
@@ 4.4 Application to Real-world PDE Dynamics
Apart from the simulated benchmark above, we also leverage two real-world PDE-governed Ocean-Atmosphere dynamics datasets, including Sea Surface Temperature (SST) (Huang et al., 2021) and stereo Sea Surface Elevation (Guimarães et al., 2020) to validate iMOOE’s generalization capability. These two datasets represent upper-ocean thermodynamics and free-surface ocean wave dynamics respectively, and contain sensory measurement noise. In SST forecasting setting, a specific region on Pacific Ocean is selected and divided into $60\times 60$ grid. We predict SST state of future $6$ days using past $4$ days observations. SST data between year 19821982-$2019$ and 20202020-$2021$ is utilized for training and testing. Note that each independent SST trajectory can be deemed as an instance from a unique environment, since daily SST variations are affected by many meteorological conditions like solar radiation and wind speed. Input parameters are also unknown so we feed one-valued vector to the fusion network. In SSE forecasting setting, we choose the wave dynamics recorded at La Jument lighthouse with total 4500 frames on $241\times 221$ grid. The training and testing sets are formed by the first 4000 frames and remaining 500 frames. We input past 4 steps SSE state to predict future 6 steps. We take a typical ocean dynamics forecasting baseline called DyAd (Wang et al., 2022b) and present OOD comparison results in Table 3, 4. We find that iMOOE can attain the lowest mean and variance on two metrics across various OOD samples, which reflects iMOOE’s capability to capture the underlying physics law in real-world ocean dynamics. Test samples in Fig. 3, 4 exhibit iMOOE can capture local variations of SST and SSE with higher fidelity and accuracy. Beyond 2D PDE-governed dynamics, we further demonstrate iMOOE’s zero-shot OOD capability can be extended to other types of dynamical systems in Appendix F.2.
@@
Number of expert $K$ DR BG Inference time
@@
4.5 Sensitivity analysis
In Table 6, we investigate the influence of the number of operator experts $K$ on iMOOE’s zero-shot OOD performance. To ensure a fair comparison, we only escalate $K$ from 1 to 4 and keep other setups like the width of FNO and training batch size unchanged. We find that the best-performing group is $K=3$ while the worst setting is $K=1$. This reveals that small $K$ (i.e. only 1 expert) is not sufficient to capture the operator invariance, while large $K$ (i.e. 4 experts) could be redundant given that actual PDE systems contains only a few number of compositional invariant operators (Rudy et al., 2017). Besides, the increasing number of neural operators can exacerbate the computational overhead. Refer to Appendix E for more detailed explanations on the effect of expert number $K$ and its distinction from the mixture-of-expert (MoE) architecture in large foundation models (LFMs). Refer to Appendix F.3 for more sensitivity analysis on loss weights in Eq. 9.
@@ 5 Conclusion
Acknowledgements
This work was supported in part by the National Key R&D Program of China (Grant No.2023YFF0725001), in part by the National Natural Science Foundation of China (Grant No.92370204), in part by the guangdong Basic and Applied Basic Research Foundation (Grant No.2023B1515120057), in part by the Key-Area Special Project of Guangdong Provincial Ordinary Universities(2024ZDZX1007).